OpenStudy (anonymous):

Bob drove 120 miles on his vacation. He drove an average of 1.2 times faster on the second 60 miles of his trip than he did on the first 60 miles of his trip. Which expression represents the time he spent driving? Let x = his speed on the first half of the trip. (1 point)

OpenStudy (anonymous):

A. 110/x B. 120/x C.132x D.132/x

OpenStudy (anonymous):

time taken to cover first 60 miles=60/x time taken to cover next 60 miles=60/1.2x=(60*10)/12x=50/x total time taken 60/x+50/x=?

OpenStudy (anonymous):

110/x ?

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thanks can you help me with 1 more its a word problem

OpenStudy (anonymous):

A custom T-shirt company charges a $200.00 set up fee and $5.00 per T-shirt. Approximately how many T-shirts must be produced so that the total cost per T-shirt is $5.33? (1 point) 1,000 t-shirts 200 t-shirts 600 t-shirts 1,266 t-shirts

OpenStudy (anonymous):

i will try to solve it.

OpenStudy (anonymous):

200+5x=5.33x 5.33x-5x=200 .33x=200 x=200/.33=606 as the rate is approximately $5 so required T-shirts=600

OpenStudy (anonymous):

okay i understand now thank you so muchhh ( :

OpenStudy (anonymous):

yw